OpenStudy (anonymous):

Solve the quadratic equation by completing the square 4x^2+24x+1=0 I get stuck when I get to the (X-3)^2= -1/4 Please help? I also know that you need to add 9 to the -1/4...I'm so confused

OpenStudy (mathstudent55):

\( 4x^2 + 24x + 1 = 0\) Divide both sides by 4: \( x^2 + 6x + \frac{1}{4} = 0 \) Subtract \( \frac{1}{4} \) from both sides: \( x^2 + 6x ~~~~~~~~~~= -\frac{1}{4} \) Add \( (\frac{b}{2})^2 \) to both sides: \( x^2 + 6x + 9= -\frac{1}{4} + 9\) \( (x + 3)^2 = -\frac{1}{4} + \frac{36}{4} \) \( (x + 3)^2 = \frac{35}{4} \)

OpenStudy (mathstudent55):

How did you get )x - 3)^2? It's (x + 3)^2.

OpenStudy (mathstudent55):

@JennyMae18 Do you follow what I did?

OpenStudy (anonymous):

Oh..I'm sorry. I mistyped. It's -24x not plus...

OpenStudy (anonymous):

It's the same steps though, right?

OpenStudy (anonymous):

\[4x ^{2}+24x+1=0\] \[x ^{2}+6x=-\frac{ 1 }{ 4 }\] adding both sides (6/2 )^2 i.e 9 \[x ^{2}+6x+9=\frac{ -1 }{4}+9=\frac{ 35 }{ 4 }\] \[\left( x+3 \right)^{2}=\left( \frac{ \sqrt{35} }{2 } \right)^{2}\] \[x+3=\pm \frac{ \sqrt{35} }{ 2 }\] \[x=-3+\frac{ \sqrt{35} }{ 2 },x=-3-\frac{ \sqrt{35} }{2}\]

OpenStudy (mathstudent55):

yes

OpenStudy (mathstudent55):

\(4x^2 - 24x + 1 = 0\) Divide both sides by 4: \( x^2 - 6x + \frac{1}{4} = 0 \) Subtract \(\frac{1}{4}\) from both sides: \( x^2 - 6x ~~~~~~~~~~= -\frac{1}{4} \) Add \( \frac{b}{2})^2\) to both sides: \( x^2 - 6x + 9= -\frac{1}{4} + 9\) \( (x - 3)^2 = -\frac{1}{4} + \frac{36}{4} \) \( (x - 3)^2 = \frac{35}{4} \)

OpenStudy (mathstudent55):

\( (x - 3)^2 = \frac{35}{4} \) \( x - 3 = \pm \sqrt{ \frac{35}{4} }\) \( x = 3 \pm \frac{ \sqrt{35} }{2} \)